how would i solve for x in this problem:
4/9(x+3)=g

Answer by chessace(471)

Do whatever legal it takes to get x by itself on the left:
Assuming it's really (4/9)(x+3) = g
Mult by recipical of left factor, i.e. 9/4:
x+3 = 9g/4
Subt 3:
x = 9g/4 -3
Assuming it really is 4/(9*(x+3)) = g
Still mult by 9/4:
1/(x+3) = 9g/4
Mult by x+3:
1 = (9g/4) * (x+3)
Flip sides and div by (9g/4):
x+3 = 4/(9g)
x = 4/(9g) - 3